# Evaluating LLMs and Pre-trained Models for Text Summarization Across Diverse Datasets

### v1

@@ V-B Case study
Ground-truth summary: Fire breaks out at the General Electric Appliance Park in Louisville, Kentucky. City officials report no one is believed to be injured or trapped.
BART: Fire damages industrial park in Kentucky, official says. Video shows blaze spreading thick smoke over area. No injuries or trapped, official says. Officials didn’t know what had caused the fire, which had gone to at least four alarms. Officials didn’t know what had caused the fire.
FLAN-T5: Fire breaks out in Kentucky industrial park, sending plumes of thick smoke over the area. No reports of anyone injured or trapped in the blaze, officials say. GE’s facility in Louisville Appliance Park is “revitalizing manufacturing in the United States.”
@@
Gemma-7B: Massive fire breaks out at General Electric Appliance Park in Louisville, Kentucky. Firefighters take up positions around affected buildings, spraying water. Authorities don’t know what caused the fire, which has gone to at least four alarms.
@@
Ground-truth summary Sri lanka closes schools as war escalates.
BART generated: Sri lanka government shuts government schools amid military campaign against rebels.
FLAN-T5 generated: Sri lanka announces closure of government schools as tamil separatists campaign intensifies in north of london.
LLaMA-3-8B generated: Srilanka shuts schools as war intensifies.
Gemma-7B generated: Sri lanka closes schools amid escalating war.

### v2

@@ V-B Case study
Ground-truth summary: “Fire breaks out at the General Electric Appliance Park in Louisville, Kentucky. City officials report no one is believed to be injured or trapped.”
BART: “Fire damages industrial park in Kentucky, official says. Video shows blaze spreading thick smoke over area. No injuries or trapped, official says. Officials didn’t know what had caused the fire, which had gone to at least four alarms. Officials didn’t know what had caused the fire.
FLAN-T5: “Fire breaks out in Kentucky industrial park, sending plumes of thick smoke over the area. No reports of anyone injured or trapped in the blaze, officials say. GE’s facility in Louisville Appliance Park is “revitalizing manufacturing in the United States”.
@@
Gemma-7B: “Massive fire breaks out at General Electric Appliance Park in Louisville, Kentucky. Firefighters take up positions around affected buildings, spraying water. Authorities don’t know what caused the fire, which has gone to at least four alarms.
@@
Ground-truth summary “Sri lanka closes schools as war escalates.”
BART generated: “Sri lanka government shuts government schools amid military campaign against rebels.”
FLAN-T5 generated: “Sri lanka announces closure of government schools as tamil separatists campaign intensifies in north of london.”
LLaMA-3-8B generated: “Srilanka shuts schools as war intensifies.”
Gemma-7B generated: “Sri lanka closes schools amid escalating war.”